OpenStudy (anonymous):

1. At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90. a. Write a system of two equations to model this problem. Let j stand for the price of a glass of juice and let p stand for the price of a pancake special. b. Use the substitution method to solve the system of equations. c. What would be the bill for one glass of orange juice and one pancake special?

OpenStudy (anonymous):

= I have no idea. So far I got $7.60 = 3OJ + P OJ

OpenStudy (anonymous):

2.90 = 1OJ + 2P

ganeshie8 (ganeshie8):

thats a good start ! lets call the price of orange juice = \(j\)

ganeshie8 (ganeshie8):

then, for partA you get : \(\large 7.6 = 3j + 5p \) \(\large 2.9 = 1j + 2p\)

ganeshie8 (ganeshie8):

see if that looks okay...

OpenStudy (anonymous):

Yea thats what I got

ganeshie8 (ganeshie8):

good, we're done wid partA.

OpenStudy (anonymous):

How do we solve the third part of the question

ganeshie8 (ganeshie8):

did u finish partB ?

OpenStudy (anonymous):

No not yet

ganeshie8 (ganeshie8):

you need to finish partB first

OpenStudy (anonymous):

Alright

ganeshie8 (ganeshie8):

isolate \(j\) from the second equation, then substitute it in the first equation

ganeshie8 (ganeshie8):

your second equation : \(\large 2.9 = 1j + 2p\) \(\large j = 2.9 - 2p\)

ganeshie8 (ganeshie8):

substitute this value in first equation : \(\large 7.6 = 3(2.9-2p) + 5p \)

ganeshie8 (ganeshie8):

solve for \(\large p\), can u ?

OpenStudy (anonymous):

7.6=8.7-6p+5p

ganeshie8 (ganeshie8):

yes ! solve it all the way

OpenStudy (anonymous):

7.6=8.7+11p?

ganeshie8 (ganeshie8):

try again whats -6 + 5 ?

OpenStudy (anonymous):

Oops -1

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

7.6=8.7-1p

ganeshie8 (ganeshie8):

Yep, isolate \(\large p\)

ganeshie8 (ganeshie8):

7,6 = 8.7 - p p = ?

OpenStudy (anonymous):

1.1-1p?

ganeshie8 (ganeshie8):

7,6 = 8.7 - p 7.6 - 8.7 = -p -1.1 = -p 1.1 = p

ganeshie8 (ganeshie8):

so the price of one pancake is $1.1

ganeshie8 (ganeshie8):

plug this value of \(p = 1.1\) in any one of the equations and solve \(j\) also

ganeshie8 (ganeshie8):

maybe plug it in second equation : \(\large 2.9 = 1j + 2p\) \(\large 2.9 = 1j + 2*1.1 \) \(\large 2.9 = j + 2.2\) \(\large 0.7 = j\) \(\large j = 0.7\)

ganeshie8 (ganeshie8):

So the solution to system of equations is : \(p = 1.1,~~j = 0.7\)

ganeshie8 (ganeshie8):

we're done wid partB

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense...

OpenStudy (anonymous):

7.6=3(2.9-2*1.1)=5*1.1

OpenStudy (anonymous):

It makes since

ganeshie8 (ganeshie8):

good, see if u can do partC

ganeshie8 (ganeshie8):

you can use the prices from partB to do partC

OpenStudy (anonymous):

$1.01 pancake $0.7 Juice?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

nooo, wait

ganeshie8 (ganeshie8):

we got : $1.1 panccale, and $0.7 juice right ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

okay use them

OpenStudy (anonymous):

Ok so thats part C

ganeshie8 (ganeshie8):

what we did so far is PartB

ganeshie8 (ganeshie8):

for PartC, you need to calculate the total price u nee to pay for one pancake and one juice

OpenStudy (anonymous):

$1.71?

ganeshie8 (ganeshie8):

thats it !!

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

Thank You for the help

ganeshie8 (ganeshie8):

you're welcome !

OpenStudy (myth1111):

I know this is from 10 months ago, but you guys messed up on the easiest part ^_^. It's not 1.01 + 0.7 = 1.71, it's 1.1 + 0.7 = 1.80, so it would be $1.80 for Part C.